james has 30cents and Tom has $1.86 . If James saves 5 cents a day and Tom spends 7 cents a day ,after how many days will they have equal amounts?
Found 4 solutions by Alan3354, ankor@dixie-net.com, Theo, MathTherapy:
Answer by Alan3354(69443)

james has 30 cents and Tom has $1.86. If James saves 5 cents a day and Tom spends 7 cents a day, after how many days will they have equal amounts?
--------
d = # of days
30 + 5d = 186 - 7d

Answer by ankor@dixie-net.com(22740)

james has 30cents and Tom has $1.86 .\ If James saves 5 cents a day and Tom spends 7 cents a day, after how many days will they have equal amounts?
;
let d = no. of days till they have equal amts
:
change 1.86 to cents
5d + 30 = 186 - 7d
5d + 7d = 186 - 30
12d = 156
d = 156/12
d = 13 days
:
:
Check, find out how much each will actually have
5(13) + 30 = 95
186 - 7(13) = 95

Answer by Theo(13342)

x = number of days.
j = number of cents that james had.
t = number of cents that tom has.
formula for james is j = 30 + 5x
formula for tom is t = 186 - 7x
everything is in cents.
they will have equal amounts when j = t
when j = t, 30 + 5x = 186 - 7x
your formula to solve is 30 + 5x = 186 - 7x
subtract 30 from both sides of this equation and add 7x to both sides of this equation to get:
5x + 7x = 186 - 30
combine like terms to get:
12x = 156
divide both sides of this equation to get:
x = 156 / 12
solve for x to get x = 13
they will have the same amount of money in 13 days.
that should be your answer.

you could have also done the same problem dealing with dollars rather than cents.
in that case:
formula for james is j = .3 + .05x
formula for tom is t = 1.86 - .07x
they will have the same amounts when j = t.
in that case, .3 + .05x = 1.86 - .07x
subtract .3 from both sides of this equation and add .07x to both sides of this equation to get:
.05x + .07x = 1.86 - .3
combine like terms to get:
.12x = 1.56
divide both sides of this equation by .12 to get:
x = 1.56 / .12
solve for x to get x = 13
same solution, just dealing with dollars rather than cents.

Answer by MathTherapy(10579)


james has 30cents and Tom has $1.86 . If James saves 5 cents a day and Tom spends 7 cents a day ,after how many days will they have equal amounts?
Let number of days be d
Then in "d" days, James will have .3 + d(.05), or .3 + .05d
Also, in "d" days, Tom will have: 1.86 - d(.07), or 1.86 - .07d
To have the same amount, we set them equal to get: .3 + .05d = 1.86 - .07d
.05d + .07d = 1.86 - .3 ------ Adding .07d, and subtracting .3 to/from both sides
.12d = 1.56
Number of days for them to have an equal amount, or: